# Reconciling Reality through Simulation: A Real-to-Sim-to-Real Approach for Robust Manipulation

### v2

@@ III-B Impact of Co-Training with Real-World Data
III-C Is Real-to-Sim Transfer Necessary?
@@
III-C1 Real-to-Sim Transfer of Scenes
@@
To probe the benefits of RialTo over such a sim-only training pipeline, we compared the performance against a policy trained using only synthetic assets. Using an amount of time effort roughly comparable to what is required from a single user following our real-to-sim approach (see Section V), we collected a set of 4 drawers from the Objaverse dataset (see Figure 6). Although this is small compared to the growing size of 3D object datasets, we found it non-trivial to transfer articulated objects into simulation-ready USDs and we leave it as future work. Given these manually constructed diverse simulation scenes, we then trained a multi-task policy using RialTo from 20 demonstrations to open the 4 drawers. See Appendix VIII-C for the minor modifications to incorporate multi-task policy learning to RialTo.
@@
III-C2 Real-to-sim transfer of policies
@@
Figure 5 shows the real-world performance of policies trained using RialTo when starting the RL fine-tuning step using real-world demonstrations as explained in II-C1 against using demonstrations provided directly in simulation. We observe that the performance for both cases is very close. These results show that RialTo successfully learns policies with demonstrations from either source of supervision as long as we keep co-training the policies with real-world data in the teacher-student distillation step. Firstly, this indicates that we do not need to collect both demos in sim and real, but we can run RialTo uniquely from the demos in the real world. Furthermore, this flexibility is a strength of our pipeline, as the ease of acquiring different sources of supervision may vary across deployment scenarios – i.e., one could use policies pretrained from large-scale real-world data or obtain data from a simulation-based crowdsourcing platform.
III-D Scaling RialTo to In-the-Wild Environments
@@
In this section, we scale up RialTo to more uncontrolled and in-the-wild environments. We test RialTo on three different tasks: open the microwave in a kitchen (also shown in Section III-A), put a cup in the trash, and bring the plate from the sink to the dishrack. We observe that RialTo scales up to these more diverse scenes and continues to perform significantly better than standard imitation learning techniques. In particular, RialTo brings on average a 57% improvement upon standard imitation learning, see Fig 7.
IV Further Analysis and Ablations
@@
IV-A Training with Distractors
@@
IV-B Comparison to RL from Scratch
@@
IV-C RL from Vision
RialTo’s “inverse distillation” procedure to a compact state-space adds some methodological overhead to the system when compared to the possibility of doing RL fine-tuning directly on visual observations. However, as reported in Appendix Fig. 13, on the task of drawer opening, RL from compact states achieves a 96% success rate after 12 hours of wall-clock time, while RL from vision only achieves a 1% success rate after 35 hours. Hence, inverse distilling to state space is necessary because training RL from vision with sparse rewards is prohibitively slow, motivating the methodology outlined in Section II-C1.
V User study
We analyzed the usability of RialTo’s pipeline for bringing real-world scenes to simulation. We ran a user study over 6 people, User 6 being an expert who used the GUI before and Users 1-5 never did any work on simulators before. Each participant was tasked with creating an articulated scene using the provided GUI. More precisely, their task was to: 1) scan a big scene, 2) cut one object, 3) scan and upload a smaller object, and 4) add a joint to the scene. From Figure 8, we found that the average total time to create a scene was 25 minutes and 12 seconds of which only 14 minutes and 40 seconds were active work. We also observed that the expert user accomplished the task faster than the rest, and twice as fast as the slowest user. This indicates that with practice, our GUI allows users to become faster at generating scenes. We conclude that doing the real-to-sim transfer of the scenes using the proposed GUI seems to be an intuitive process that is neither time nor labor-intensive when compared to collecting many demonstrations in the real world. We provide more details about the study in Appendix XII.
@@ VI Limitations and Conclusion
While our use of 3D point clouds instead of RGB enables easier sim-to-real transfer, we require accurate depth sensors that can struggle to detect thin, transparent, and reflective objects. Future work may investigate applying RialTo to train policies that operate on RGB images or RGBD, as our framework makes no fundamental assumptions that prevent using different sensor modalities. We are also limited to training policies for tasks that can be easily simulated and for real-world objects that can be turned into digital assets. Currently, this is primarily limited to articulated rigid bodies, but advancements in simulating and representing deformables should allow our approach to be applied to more challenging objects. Even though we show RialTo works on fast controllers, these are still relatively slow to minimize the sim-to-real gap in dynamics, thereafter there is potential to investigate tasks for which faster controllers are needed. In this work, we consider relatively quasistatic problems, where exact identification of physics parameters is not necessary for the constructed simulation. This will become important as more complex environments are encountered. Finally, as we explain in Section XIII, RialTo currently takes around 2 days of wall-clock time end-to-end to train a policy for each task, this time bottleneck makes continual learning infeasible and understanding how to obtain policies faster with minimal human supervision would be valuable. We expect with more efficient techniques for learning with point clouds and better parallelization, this procedure can be sped up significantly.

### v3

@@ IV-B Impact of Co-Training with Real-World Data
IV-C Is Real-to-Sim Transfer Necessary?
@@
IV-C1 Real-to-Sim Transfer of Scenes
@@
To probe the benefits of RialTo over such a sim-only training pipeline, we compared the performance against a policy trained using only synthetic assets. Using an amount of time effort roughly comparable to what is required from a single user following our real-to-sim approach (see Section VI), we collected a set of 4 drawers from the Objaverse dataset (see Figure 7). Although this is small compared to the growing size of 3D object datasets, we found it non-trivial to transfer articulated objects into simulation-ready USDs and we leave it as future work. Given these manually constructed diverse simulation scenes, we then trained a multi-task policy using RialTo from 20 demonstrations to open the 4 drawers. See Appendix IX-C for the minor modifications to incorporate multi-task policy learning to RialTo.
@@
IV-C2 Real-to-sim transfer of policies
@@
Figure 6 shows the real-world performance of policies trained using RialTo when starting the RL fine-tuning step using real-world demonstrations as explained in III-C1 against using demonstrations provided directly in simulation. We observe that the performance for both cases is very close. These results show that RialTo successfully learns policies with demonstrations from either source of supervision as long as we keep co-training the policies with real-world data in the teacher-student distillation step. Firstly, this indicates that we do not need to collect both demos in sim and real, but we can run RialTo uniquely from the demos in the real world. Furthermore, this flexibility is a strength of our pipeline, as the ease of acquiring different sources of supervision may vary across deployment scenarios – i.e., one could use policies pretrained from large-scale real-world data or obtain data from a simulation-based crowdsourcing platform.
IV-D Scaling RialTo to In-the-Wild Environments
@@
In this section, we scale up RialTo to more uncontrolled and in-the-wild environments. We test RialTo on three different tasks: open the microwave in a kitchen (also shown in Section IV-A), put a cup in the trash, and bring the plate from the sink to the dishrack. We observe that RialTo scales up to these more diverse scenes and continues to perform significantly better than standard imitation learning techniques. In particular, RialTo brings on average a 57% improvement upon standard imitation learning, see Fig 8.
V Further Analysis and Ablations
@@
V-A Training with Distractors
@@
V-B Comparison to RL from Scratch
@@
V-C RL from Vision
RialTo’s “inverse distillation” procedure to a compact state-space adds some methodological overhead to the system when compared to the possibility of doing RL fine-tuning directly on visual observations. However, as reported in Appendix Fig. 14, on the task of drawer opening, RL from compact states achieves a 96% success rate after 12 hours of wall-clock time, while RL from vision only achieves a 1% success rate after 35 hours. Hence, inverse distilling to state space is necessary because training RL from vision with sparse rewards is prohibitively slow, motivating the methodology outlined in Section III-C1.
VI User study
We analyzed the usability of RialTo’s pipeline for bringing real-world scenes to simulation. We ran a user study over 6 people, User 6 being an expert who used the GUI before and Users 1-5 never did any work on simulators before. Each participant was tasked with creating an articulated scene using the provided GUI. More precisely, their task was to: 1) scan a big scene, 2) cut one object, 3) scan and upload a smaller object, and 4) add a joint to the scene. From Figure 9, we found that the average total time to create a scene was 25 minutes and 12 seconds of which only 14 minutes and 40 seconds were active work. We also observed that the expert user accomplished the task faster than the rest, and twice as fast as the slowest user. This indicates that with practice, our GUI allows users to become faster at generating scenes. We conclude that doing the real-to-sim transfer of the scenes using the proposed GUI seems to be an intuitive process that is neither time nor labor-intensive when compared to collecting many demonstrations in the real world. We provide more details about the study in Appendix XIII.
@@ VII Limitations and Conclusion
While our use of 3D point clouds instead of RGB enables easier sim-to-real transfer, we require accurate depth sensors that can struggle to detect thin, transparent, and reflective objects. Future work may investigate applying RialTo to train policies that operate on RGB images or RGBD, as our framework makes no fundamental assumptions that prevent using different sensor modalities. We are also limited to training policies for tasks that can be easily simulated and for real-world objects that can be turned into digital assets. Currently, this is primarily limited to articulated rigid bodies, but advancements in simulating and representing deformables should allow our approach to be applied to more challenging objects. Even though we show RialTo works on fast controllers, these are still relatively slow to minimize the sim-to-real gap in dynamics, thereafter there is potential to investigate tasks for which faster controllers are needed. In this work, we consider relatively quasistatic problems, where exact identification of physics parameters is not necessary for the constructed simulation. This will become important as more complex environments are encountered. Finally, as we explain in Section XIV, RialTo currently takes around 2 days of wall-clock time end-to-end to train a policy for each task, this time bottleneck makes continual learning infeasible and understanding how to obtain policies faster with minimal human supervision would be valuable. We expect with more efficient techniques for learning with point clouds and better parallelization, this procedure can be sped up significantly.